Which ordered pair is a solution of the equation
-4x+7=2y-3

Found 3 solutions by Alan3354, addingup, MathTherapy:
Answer by Alan3354(69443)

Which ordered pair is a solution of the equation
-4x+7=2y-3
-----------------
There are an infinite number of them.

Answer by addingup(3677)

-4x+7 = 2y-3 reverse the equation
2y-3 = -4x+7
2y = -4x+10
y = -2x+5 now we have the equation in slope-intercept form y = mx+b.
The equation tells us that the slope m = -2 but since slope = rise/run, we need to have a fraction here. -2 as a fraction is -2/1, so our rise is -2 and our run is 1.
The intercept, the point at which y crosses over, is 5. I'm going to plot this to get your ordered pair:

Answer by MathTherapy(10551)


Which ordered pair is a solution of the equation
-4x+7=2y-3
I don't know why you're asking for help, you state "which pair" and there are NO PAIRS to choose from. What's the sense in asking for help yet you do not provide the entire problem?
I'll give you a HINT: (5, 0) and (- 2, 1) ARE NOT solution-pairs, as a person who responded has stated.
Lesson learnt: Never trust some people on this site to give you correct answers. Some are wrong sometimes, some MOST times, some almost never, or seldom, and a few are,
probably 99% of the times they respond.
BEWARE!!!! Don't say I didn't warn you!